### v1

# Temporal Evolution of the Third Interstellar Comet 3I/ATLAS: Spin, Color, Spectra and Dust Activity††thanks: based on observations made with the Southern African Large Telescope (SALT)

Abstract
Aims. We aim to characterize the physical and activity properties of the interstellar comet 3I/ATLAS through spectroscopic and photometric observations during the first month after its discovery.
Methods. We performed time-series photometry and long-slit spectroscopy between 2 and 29 July 2025 using multiple ground-based telescopes. Photometric data were calibrated against field stars from the ATLAS and APASS catalogs, and Fourier analysis was applied to derive the comet’s rotational period. Spectral data were obtained using the SALT telescope and Nordic Optical Telescope.
Results. We report a spin period of h with a lightcurve amplitude of approximately 0.3 mag. The comet exhibits increasing dust activity and reddening colors during the observation period, with no visible tail detected, likely due to viewing geometry and low dust production. Dust mass loss rates are estimated between 0.3 and 4.2 kg s-1, consistent with weakly active distant comets. Spectral colors are similar to those of outer Solar System comets and differ from previously reported values for 3I/ATLAS.
Conclusions. The morphological and photometric properties of 3I/ATLAS are consistent with a weakly active comet of outer Solar System origin, despite its interstellar provenance. Continued monitoring around perihelion is necessary to track changes in activity, color, which will provide insights into the evolution of interstellar materials under solar radiation.
Key Words.:
comets — interstellar objects — 3I/ATLAS — techniques: photometric — techniques: spectroscopic — methods: observational1 Introduction
Interstellar objects (ISOs) are minor bodies on clearly hyperbolic, Sun-unbound trajectories; their excess speed at infinity confirms an origin beyond the Solar System (Meech et al., 2017). Each ISO offers a physical sample of extrasolar planetesimal material. Studying ISOs is compelling for several reasons: their compositions, activity, and colors probe solid-body formation in other planetary systems; the contrast between the first two ISOs—1I/‘Oumuamua, an apparently inactive, extremely flattened body, and 2I/Borisov, a dust-rich comet—demonstrates the diversity of extrasolar debris reservoirs (Micheli et al., 2018; Jewitt & Luu, 2019); their trajectories reveal how Galactic tides, stellar flybys, and giant-planet scattering shape Oort clouds (Seligman & Moro-Mart´ın, 2023); and their unusual orbits offer rare tests of sublimation and space weathering. Rapid-response ISO campaigns also refine observing strategies ahead of expected ISO discoveries by the Rubin Observatory (Fitzsimmons et al., 2023).
The third confirmed ISO, comet 3I/ATLAS (C/2025 N1), was discovered on 1 July 2025 by the southern ATLAS 0.5 m telescope at Río Hurtado, Chile, and announced as interstellar the next day (Minor Planet Center, 2025). Archival images from ZTF and ATLAS extended the orbit back to mid-June and confirmed an extreme eccentricity, (Seligman et al., 2025). Early follow-up showed that 3I/ATLAS was already active at au, with a dust coma detected by multiple facilities (Bolin et al., 2025; de la Fuente Marcos et al., 2025).
According to JPL Horizons (Giorgini, 2025), 3I/ATLAS will reach perihelion on 29–30 October 2025 at au (just inside Mars’ orbit), then recede into interstellar space. It becomes unobservable from the ground in September due to low solar elongation but re-emerges in the northern sky in late November. Models suggest it will remain visible to ground-based and space telescopes into mid-2026.
In this paper, we present optical photometry and low-resolution spectroscopy of 3I/ATLAS obtained during its first month of observation, while still beyond 3.5 au, establishing a pre-perihelion baseline for tracking its evolution.
2 Observations
2.1 Spectroscopy
Low-resolution spectroscopic observations of 3I/ATLAS (henceforth 3I) were conducted on 15 and 25 July using the 10-meter Southern African Large Telescope (SALT; see Section A.2.1) and the 2.56-meter Nordic Optical Telescope (NOT; see Section A.2.2), respectively. The SALT spectrum covered the visible wavelength range from to , while for the NOT spectrum we used data from to . To remove the solar continuum and derive the reflectance spectrum of the comet, we observed the solar analogue star SA112-1333 (Landolt, 1992) on the same nights. Spectra of the spectrophotometric standard star Feige 110 (Oke, 1990) were also obtained for flux calibration. Standard data reduction procedures were applied. Individual comet spectra were median-combined and then divided by the solar analogue spectrum. The resulting reflectance spectra are shown in Fig. 1.
2.2 Photometry
Photometric observations of 3I were conducted between 2 and 29 July 2025 using the Faulkes Telescope North and South (FTN and FTS; see Section A.1.1), the Nordic Optical Telescope (NOT; see Section A.1.2), the Telescopi Joan Oró (TJO; see Section A.1.3), the Lesedi Telescope (see Section A.1.4), and the Skalnaté Pleso Observatory (see Section A.1.5). Time-series photometry was performed exclusively in the SDSS band, while multi-band imaging in and VR filters were used for color measurements. Exposure times were adjusted dynamically based on the seeing conditions and plate scale, maximizing integration time while ensuring stellar point sources remained untrailed. All images were calibrated using standard procedures, including bias subtraction, dark-frame correction, and flat-fielding. Photometric measurements were obtained through aperture photometry relative to selected field stars. Calibration was performed using reference magnitudes from the ATLAS All-Sky Stellar Reference Catalog (Tonry et al., 2018) or from the APASS catalog (Henden & Munari, 2014).
3 Results
3.1 Analysis of the color evolution
In the range between – we observed a slight trend of reddening in the comet’s color as it approaches the Sun (see rising and colors in Fig 4), a neutral evolution in the range – (see constant color in Fig 4), and a slight bluing beyond (see decrease in color in Fig 4). However, in general, the measured color remains bluer than reported by Bolin et al. (2025) for 3I at comparable heliocentric distances but our estimates are more consistent with those of Belyakov et al. (2025) and Opitom et al. (2025), who also found relatively neutral to slightly red colors for 3I. When compared to other objects, the color of comet 3I closely resembles that reported for 2I/Borisov (Epifani et al., 2021) at heliocentric distances near 2 au. Moreover, it is similar to distant comets observed at comparable distances around 4 au (e.g., Snodgrass et al. 2008; Ivanova et al. 2015; Voitko et al. 2024), suggesting analogous dust properties among these bodies.
A similar trend is evident in the spectroscopic data, where we derived spectral gradient values from normalized reflectance measurements obtained with NOT and SALT. Based on observations from 4 July, Seligman et al. (2025) reported a gradient of . Our analysis shows an increase to from the SALT data on 15 July, and from NOT observations on 25 July, both calculated over the – range. This progressive reddening over time suggests an evolving surface or coma composition. Beyond , the NOT spectrum becomes noticeably flatter, trending towards a more neutral color. Compared to earlier results by Opitom et al. 2025 and de la Fuente Marcos et al. 2025, we observe a decrease in spectral slope to within –. Future follow-up spectroscopy is planned to track the evolution of the spectral properties, including potential emission lines and ongoing activity indicators.
3.2 Morphology of the coma
We created intensity maps (Fig. 2) by co-adding the photometric images taken with the SDSS and filters on the two dates of the observing campaign, 4 and 29 July 2025. A more detailed view of the comet’s morphological evolution from 2 to 29 July is presented in Fig. 5. The cometary coma appears compact and slightly asymmetrical. No visible tail was detected before 16 July. A slight elongation of the coma, resembling a short tail, was observed between 18 and 26 July, directed roughly at a position angle of . Observations by Kareta et al. (2025) using the NASA Infrared Telescope Facility similarly showed no prominent tail. This absence, also evident in our images and those from other observers, can be partially attributed to the tail being projected ‘behind’ the comet due to the low phase angle during the observations. This effect may be further enhanced if the comet contains relatively few fine dust grains, which are the most affected by radiation pressure. To search for faint structures within the weak coma of comet 3I, we applied digital image enhancement techniques, including the rotational gradient method (Larson & Sekanina, 1984), the 1/ profile, and renormalization filtering (Samarasinha & Larson, 2014). These methods were applied only to the stacked images, as the object was too faint for filtering of individual frames. No jet-like or other morphological features were detected.
3.3 Magnitude and dust production level
Between 2 and 22 July, photometric observations of 3I revealed a progressive increase in brightness, consistent with its decreasing heliocentric distance from 4.4 au to approximately 3.8 au. The absolute magnitude was determined to be approximately mag, in agreement with the value reported by Jewitt & Luu (2025). Assuming a geometric albedo , our estimate of the comet radius upper limit is about 11 km (see Table 2).
The dust activity level, expressed as , is approximately 300 cm, consistent with estimates by Bolin et al. (2025), but higher than values observed for 2I/Borisov (Guzik et al. 2020; Opitom et al. 2019). This value is comparable to those measured for Jupiter-family comets at large heliocentric distances, as well as for some distant comets exhibiting similar dust activity (Ivanova et al. 2015; Voitko et al. 2024).
Dust mass loss rates are estimated to range from roughly 0.3 to 4.2 kg s-1, slightly higher than those reported by Bolin et al. (2025), but lower than typical rates for Jupiter-family comets at large distances (Gillan et al., 2024), and comparable to distant comets observed at even greater heliocentric distances (Ivanova et al., 2015).
Detailed methodology, equations, and full tables are presented in Appendix B.
3.4 Lightcurve
Using the time-series photometry described in Section 2.2, we generated a rotational lightcurve of 3I. The comet was observed while crossing the Galactic plane, necessitating mitigation techniques to minimize stellar contamination. We applied Fourier analysis to search for the best-fitting spin period, which was determined to be h (see Figure 3), with a peak-to-peak amplitude of approximately 0.3 mag. The analysis also revealed two prominent alias periods near 8 and 24 hours (see Figure 6).
In some segments of the lightcurve—particularly those corresponding to data from F65 and Z23—the rotational modulation was not clearly detected. These observations often showed a relatively flat lightcurve, lacking distinct maxima or minima and instead clustering around an average magnitude. This behavior is likely due to a combination of observing conditions (e.g., atmospheric seeing, telescope resolution) and the comet’s intrinsic activity, which can enhance the surrounding coma and dilute the nucleus’s photometric variability. In particular, the extensive dataset from Z23 was probably acquired under such conditions, where the rotational signature was partially or entirely suppressed. The inclusion of these flat segments reduced the contrast of the true rotational signal and introduced ambiguity in the Fourier analysis, likely contributing to the emergence of alias solutions.
4 Conclusions
The morphology of comet 3I during one month of observations reveals clear signs of activity, with continuous dust emission evidenced by the increasing coma size as the heliocentric distance decreased. The absence of a visible tail is likely due to the observing geometry. The coma remained predominantly asymmetric, with no signs of fragmentation or rapid morphological changes. Overall, 3I exhibits characteristics typical of weakly active outer Solar System comets, despite its interstellar origin. Our time-series photometry indicates a spin period of h, with a maximum lightcurve amplitude of 0.3 mag measured at the start of the campaign. This amplitude decreased to 0.2 mag over time, likely as a result of increasing cometary activity, which progressively masked the rotational signal of the nucleus. The presence of enhanced activity is further supported by a reddening trend observed in the comet’s colors throughout the observation period.
We plan to continue photometric and spectroscopic monitoring, and to incorporate polarimetric observations to detect significant changes in color and light-scattering properties of the dust coma in both pre- and post-perihelion data. These observations are essential for characterizing the physical properties of interstellar material, understanding its evolution under solar radiation, and enabling meaningful comparisons with small bodies in the Solar System.
Acknowledgements.
T.S-R. acknowledges funding from Ministerio de Ciencia e Innovación (Spanish Government), PGC2021, PID2021-125883NB-C21. This work was (partially) supported by the Spanish MICIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” by the “European Union” through grant PID2021-122842OB-C21, and the Institute of Cosmos Sciences University of Barcelona (ICCUB, Unidad de Excelencia ‘María de Maeztu’) through grant CEX2019-000918-M. D.O. was supported by grant No. 2022/45/B/ST9/00267 from the National Science Centre, Poland. O.I. and M.H. thank the Slovak Grants by the Agency for Science, VEGA, grant No. 2/0059/22. This work is based on observations made at the South African Astronomical Observatory (SAAO), which is financially supported by the South African National Research Foundation (NRF). Some of the observations reported in this paper were obtained with the Southern African Large Telescope (SALT). Polish participation in SALT is funded by grant No. MEiN nr 2021/WK/01. The Joan Oró Telescope (TJO) at the Montsec Observatory (OdM) is owned by the Catalan Government and operated by the Institute of Space Studies of Catalonia (IEEC). Partly based on observations made with the Nordic Optical Telescope. The NOT data were obtained under program ID 71-411. Work of Skalnaté Pleso has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. ATLAS is primarily funded to search for near-Earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, and the South African Astronomical Observatory. This paper is based on observations made with the MuSCAT instruments, developed by the Astrobiology Center (ABC) in Japan, the University of Tokyo, and Las Cumbres Observatory (LCOGT). MuSCAT3 was developed with financial support by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), and is located at the Faulkes Telescope North on Maui, HI (USA), operated by LCOGT. MuSCAT4 was developed with financial support provided by the Heising-Simons Foundation (grant 2022-3611), JST grant number JPMJCR1761, and the ABC in Japan, and is located at the Faulkes Telescope South at Siding Spring Observatory (Australia), operated by LCOGT. Some observations were obtained by the Comet Chasers schools outreach program (https://www.cometchasers.org/), which is funded by the UK Science and Technology Facilities Council (via the DeepSpace2DeepImpact Project), the Open University, and Cardiff University. It accesses the LCOGT telescopes through the Schools Observatory/Faulkes Telescope Project (TSO2025A-00 DFET—The Schools’ Observatory), which is partly funded by the Dill Faulkes Educational Trust, and the LCO Global Sky Partners Programme (LCOEPO2023B-013). Schools making observations included The Coopers Company & Coborn School, Upminster, UK, Ysgol Gyfun Gymraeg Bro Edern, Cardiff, UK, St Marys Catholic Primary School, Bridgend, UK, Institut d’Alcarràs, Catalonia, Spain, Louis Cruis Astronomy Club and Srednja škola Jelkovec (Jelkovec High School), Zagreb, Croatia.References
- A’Hearn et al. (1984) A’Hearn, M. F., Schleicher, D. G., Millis, R. L., Feldman, P. D., & Thompson, D. T. 1984, AJ, 89, 579
- Belyakov et al. (2025) Belyakov, M., Bolin, B. T., Fremling, C., & Graham, M. 2025, The Astronomer’s Telegram, 17276, 1
- Bolin et al. (2025) Bolin, B. T., Belyakov, M., Fremling, C., & et al. 2025, arXiv e-prints [2507.05252]
- Bowell et al. (1989) Bowell, E., Hapke, B., Domingue, D., et al. 1989, in Asteroids II, ed. R. P. Binzel, T. Gehrels, & M. S. Matthews (University of Arizona Press), 524–556
- Burgh et al. (2003) Burgh, E. B., Nordsieck, K. H., Kobulnicky, H. A., et al. 2003, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4841, Instrument Design and Performance for Optical/Infrared Ground-based Telescopes, ed. M. Iye & A. F. M. Moorwood, 1463–1471
- Colomé et al. (2010) Colomé, J., Ribas, I., Francisco, X., et al. 2010, Advances in Astronomy, 2010, 183016
- de la Fuente Marcos et al. (2025) de la Fuente Marcos, R., Licandro, J., Alarcón, M. R., & et al. 2025, Astronomy & Astrophysics Letters [2507.12922], submitted; arXiv:2507.12922
- DeMeo et al. (2009) DeMeo, F. E., Binzel, R. P., Slivan, S. M., & Bus, S. J. 2009, Icarus, 202, 160
- Epifani et al. (2021) Epifani, E. M., Dotto, E., Perna, D., et al. 2021, Planetary and Space Science, 208, 105341
- Erasmus et al. (2025) Erasmus, N., Potter, S. B., van Gend, C. H. D., et al. 2025, in Revista Mexicana de Astronomía y Astrofísica, Serie de Conferencias (RMxAC), Vol. 59 (Instituto de Astronomía, Universidad Nacional Autónoma de México), 201–207
- Erasmus et al. (2024) Erasmus, N., Potter, S. B., van Gend, C. H. D. R., et al. 2024, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 13096, Ground-based and Airborne Instrumentation for Astronomy X, ed. J. J. Bryant, K. Motohara, & J. R. D. Vernet, 130968K
- Erasmus et al. (2024) Erasmus, N., Steele, I., Piascik, A. S., et al. 2024, Journal of Astronomical Telescopes, Instruments, and Systems, 10, 025005
- Fink & Rubin (2012) Fink, U. & Rubin, M. 2012, Icarus, 221, 721
- Fitzsimmons et al. (2023) Fitzsimmons, A., Bannister, M. T., & Moro-Martín, A. 2023, Annual Review of Astronomy and Astrophysics, 61, 155
- Fukui et al. (2022) Fukui, A., Narita, N., Kusakabe, N., & et al. 2022, Publications of the Astronomical Society of Japan, 74, 1361
- Gillan et al. (2024) Gillan, A. F., Fitzsimmons, A., Denneau, L., et al. 2024, The Planetary Science Journal, 5, 25
- Giorgini (2025) Giorgini, J. D. 2025, JPL HORIZONS system: Ephemeris for 3I/ATLAS, https://ssd.jpl.nasa.gov/horizons, accessed 2025 July 4
- Guzik et al. (2020) Guzik, P., Drahus, M., Rusek, K., et al. 2020, Nature Astronomy, 4, 53
- Henden & Munari (2014) Henden, A. & Munari, U. 2014, Contrib. Astron. Obs. Skalnate Pleso, 43, 518
- Henden et al. (2016) Henden, A. A., Templeton, M., Terrell, D., et al. 2016, VizieR Online Data Catalog, II
- Hsieh et al. (2021) Hsieh, H. H., Chandler, C. O., Denneau, L., et al. 2021, The Astrophysical Journal Letters, 922, L9
- Ivanova et al. (2015) Ivanova, O., Neslušan, L., Krišandová, Z. S., et al. 2015, Icarus, 258, 28
- Ivanova et al. (2016) Ivanova, O. V., Luk, I. V., Kiselev, N. N., et al. 2016, Planetary and Space Science, 121, 10
- Jewitt & Luu (2019) Jewitt, D. & Luu, J. 2019, Astrophysical Journal Letters, 886, L29
- Jewitt & Luu (2025) Jewitt, D. & Luu, J. 2025, The Astronomer’s Telegram, 17263, 1
- Jewitt & Meech (1987) Jewitt, D. & Meech, K. J. 1987, Astrophysical Journal, Part 1 (ISSN 0004-637X), vol. 317, June 15, 1987, p. 992-1001. NASA-supported research., 317, 992
- Jewitt et al. (2016) Jewitt, D., Mutchler, M., Weaver, H., et al. 2016, The Astrophysical Journal Letters, 829, L8
- Kareta et al. (2025) Kareta, T., Champagne, C., McClure, L., et al. 2025, arXiv preprint arXiv:2507.12234
- Kobulnicky et al. (2003) Kobulnicky, H. A., Nordsieck, K. H., Burgh, E. B., et al. 2003, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4841, Instrument Design and Performance for Optical/Infrared Ground-based Telescopes, ed. M. Iye & A. F. M. Moorwood, 1634–1644
- Kostov & Bonev (2018) Kostov, A. & Bonev, T. 2018, Bulgarian Astronomical Journal, 28, 3
- Kwiatkowski et al. (2009) Kwiatkowski, T., Kryszczyńska, A., Polińska, M., et al. 2009, A&A, 495, 967
- Landolt (1992) Landolt, A. U. 1992, AJ, 104, 340
- Larson & Sekanina (1984) Larson, S. & Sekanina, Z. 1984, AJ, 89, 571
- Meech et al. (2017) Meech, K. J., Weryk, R., Micheli, M., & et al. 2017, Nature, 552, 378
- Micheli et al. (2018) Micheli, M., Farnocchia, D., Meech, K. J., & et al. 2018, Nature, 559, 223
- Minor Planet Center (2025) Minor Planet Center. 2025, C/2025 N1 (ATLAS) = 3I/ATLAS, Minor Planet Electronic Circular 2025-N12, https://minorplanetcenter.net/mpec/K25/K25N12.html
- Narita et al. (2020) Narita, N., Fukui, A., Kusakabe, N., & et al. 2020, Journal of Astronomical Telescopes, Instruments, and Systems, 6, 045002
- Oke (1990) Oke, J. 1990, Astronomical Journal (ISSN 0004-6256), vol. 99, May 1990, p. 1621-1631. Research supported by the Space Telescope Science Institute., 99, 1621
- Opitom et al. (2019) Opitom, C., Fitzsimmons, A., Jehin, E., et al. 2019, Astronomy & Astrophysics, 631, L8
- Opitom et al. (2025) Opitom, C., Snodgrass, C., Jehin, E., et al. 2025, arXiv preprint arXiv:2507.05226
- Parrott (2020) Parrott, D. 2020, The Journal of the American Association of Variable Star Observers, 48, 262
- Potter et al. (2024) Potter, S. B., Erasmus, N., van Gend, C. H. D. R., et al. 2024, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 13098, Observatory Operations: Strategies, Processes, and Systems X, ed. C. R. Benn, A. Chrysostomou, & L. J. Storrie-Lombardi, 130980Y
- Samarasinha & Larson (2014) Samarasinha, N. H. & Larson, S. M. 2014, Icarus, 239, 168
- Schleicher & Bair (2011) Schleicher, D. G. & Bair, A. N. 2011, The Astronomical Journal, 141, 177
- Seligman et al. (2025) Seligman, D. Z., Micheli, M., Farnocchia, D., & et al. 2025, arXiv e-prints [2507.02757]
- Seligman & Moro-Mart´ın (2023) Seligman, D. Z. & Moro-Martín, A. 2023, Nature Astronomy, 7, 1105
- Snodgrass et al. (2008) Snodgrass, C., Lowry, S. C., & Fitzsimmons, A. 2008, Monthly Notices of the Royal Astronomical Society, 385, 737
- Tonry et al. (2018) Tonry, J. L., Denneau, L., Flewelling, H., et al. 2018, ApJ, 867, 105
- Voitko et al. (2024) Voitko, A., Zubko, E., Ivanova, O., Husárik, M., & Videen, G. 2024, Icarus, 411, 115967
- Willmer (2018) Willmer, C. N. A. 2018, ApJS, 236, 47
- Worters et al. (2016) Worters, H. L., O’Connor, J. E., Carter, D. B., et al. 2016, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9908, Ground-based and Airborne Instrumentation for Astronomy VI, ed. C. J. Evans, L. Simard, & H. Takami, 99083Y
Appendix A Observational facilities
A.1 Photometry
A.1.1 Faulkes Telescopes North and South
We used the 2.0-m Faulkes Telescope North (FTN; MPC code F65), located at Haleakalā Observatory in Maui, Hawai‘i (USA), and the 2.0-m Faulkes Telescope South (FTS; MPC code E10), located at Siding Spring Observatory in New South Wales (Australia), to obtain photometry in the SDSS , , , and bands using the MuSCAT3 and MuSCAT4 instruments, respectively. MuSCAT3 is equipped with four 1k 1k CCDs (Hamamatsu Photonics), each with a pixel scale of 0.27″ and a field of view of 7.4 7.4 arcmin2, while MuSCAT4 features four 2k 2k CCDs, with a pixel scale of 0.38″ and a field of view of 13 13 arcmin2 (Narita et al. 2020; Fukui et al. 2022). Both instruments enable simultaneous multi-band imaging via dichroic beam splitters. The detectors exhibit low readout noise (3–5 e-), and the photometric data were processed using the MuSCAT pipeline, which includes standard calibrations (bias, dark, and flat-field corrections) and aperture photometry optimized for each band.
A.1.2 Nordic Optical Telescope
Photometric observations in the Sloan filter were performed at the Nordic Optical Telescope (NOT) using the Alhambra Faint Object Spectrograph and Camera (ALFOSC). The NOT is a 2.56-meter telescope located at the Spanish “Roque de los Muchachos” Observatory (ORM), La Palma, Canarias, Spain (MPC code: 950). On 17 July 2025, three series of 80 exposures with 30-second integration times were obtained, and on 18 July 2025, one series of 80 exposures with 30-second integration times was acquired.
A.1.3 Telescopi Joan Oró
We used the 0.8-m Telescopi Joan Oró (TJO; MPC code C65), located in the Catalan Pyrenees, to obtain time-series photometry in the SDSS band on the nights of 4, 6, 13, 16, 17, and 19 July. On 4 and 6 July, the backup camera MEIA2 (2k 2k Andor iKon XL, pixel scale 0.36″) was used, while on the remaining nights the main camera MEIA3 (4k 4k Andor iKon XL 230-84, pixel scale 0.4″) was employed. Both cameras have very similar performance, differing slightly in gain (1.55 e-/ADU for MEIA2 vs. 1.04 e-/ADU for MEIA3) and readout noise (8 e- vs. 9 e-, respectively). The main distinction lies in the field of view: 0.13 deg for MEIA2 and 0.45 deg for MEIA3. All images were calibrated using the IEEC Calibration and Analysis Tool (ICAT; Colomé et al. 2010).
A.1.4 Lesedi
Multi-filter photometric observations were collected with the f/8, 1-meter Lesedi telescope (Worters et al. 2016), located at the South African Astronomical Observatory’s (SAAO) observing site in the Northern Cape, South Africa (GPS coordinates: 32°22′47.2′′ S, 20°48′38.4′′ E; MPC code: M28). The instrument used was Mookodi (Erasmus et al. 2024), which in imaging mode provides a field of view of 10 arcmin 10 arcmin and is equipped with a full SDSS filter set (). The telescope operates in a fully robotic mode each night as part of the SAAO’s “Intelligent Observatory” (IO) initiative (Potter et al. 2024; Erasmus et al. 2024, 2025), and data were collected under the rapid-response near-Earth asteroid follow-up program currently running at the SAAO (PIs: T.S Ngwane and N. Erasmus).
Observations were obtained on 02, 07, 15 and 23 July 2025 using 11 binning (0.59 arcsec/pixel), the slowest readout mode of 0.5 MHz (to minimize readout noise), and the highest gain setting of 1.2 e-/ADU.
On 02/07, a time-series observation was carried out by cycling through the , , , and filters continuously between 20:06:02 and 20:45:05, with an exposure time of 120 seconds per filter. The five -filter and five -filter frames collected during this window were median-stacked and used to generate the intensity maps shown in the top row of Fig. 5.
On 07/07, two time-series datasets were obtained. The first was in the filter only, spanning approximately one hour between 19:35:27 and 20:20:59, using 30-second exposures. The second followed a similar multi-filter sequence to that of 02/07, with 30-second exposures between 20:26:18 and 21:15:12. The seven -filter and six -filter frames collected during this window were median-stacked.
On 15/07, a time-series cycling between only the and filters was performed, with 30-second exposures between 19:48:07 and 20:42:47. Approximately 20 frames per filter were collected and stacked to generate the intensity maps shown in the third row of Fig. 5.
A near-identical observing strategy to that used on 15/07 was employed on 23/07, and the resulting analysis is shown in the fifth row of Fig. 5.
A.1.5 Skalnaté Pleso Observatory
Photometric material at the Skalnaté Pleso Observatory (MPC code: 056) for the interstellar comet 3I/ATLAS was acquired using a 0.61-m f/4.3 Newtonian telescope equipped with an SBIG ST-10XME CCD camera and Johnson-Cousins and photometric filters. Observations were conducted using binning, yielding a resolution of 1.07 arcsec per pixel. Standard reduction procedures were applied, including bias subtraction, dark frame subtraction, and flat-field correction. Tycho Tracker software (Parrott 2020) was used as the main tool for deriving magnitudes and quantities. Stellar magnitudes for standard stars were taken from the ATLAS catalog, an all-sky reference catalog containing approximately one billion stars down to a limiting magnitude of 19. The catalog integrates data from Pan-STARRS DR1, ATLAS Pathfinder, ATLAS re-flattened APASS, SkyMapper DR1, APASS DR9, Tycho-2, and the Yale Bright Star Catalog. Developed by Tonry et al. (2018), it provides a robust foundation for photometric calibration (Kostov & Bonev 2018). Interstellar comet 3I/ATLAS was observed on five nights between July 3 and July 23.
A.2 Spectroscopy
A.2.1 Southern African Large Telescope
Low-resolution spectroscopy of 3I/ATLAS was performed using the 10-meter Southern African Large Telescope (SALT; Kwiatkowski et al. 2009), located in Sutherland, South Africa (IAU code: B31). Observations were conducted on 2025 July 15 using the Robert Stobie Spectrograph (RSS; Burgh et al. 2003; Kobulnicky et al. 2003) under program 2025-1-DDT-004 (PI: T. Kwiatkowski). The RSS is equipped with a detector comprising a mini-mosaic of CCD chips, each with pixels of 15 m size and a pixel scale of 0.12 arcseconds per unbinned pixel. We obtained 20 exposures of 120 s each using the PG0700 grating and a narrow slit, aligned along the comet’s trail. Observations were conducted in binning mode.
Due to the comet’s proximity to the Milky Way, significant contamination from background stars occurred along the slit. Consequently, only 5120 s exposures, during which the comet was positioned against a clean sky region, were retained. All spectra obtained along the slit were extracted, but only the comet spectrum was analyzed. A solar analogue star, SA112-1333, was observed on the same night and at a similar airmass for calibration purposes.
Data reduction and calibration followed standard procedures, including bias subtraction, flat-field correction, sky background subtraction, and one-dimensional aperture extraction. Wavelength calibration was performed using an argon lamp spectrum. The individual spectra of 3I/ATLAS were median-combined and divided by the median-combined spectrum of the solar analogue to remove the solar continuum. The final reflectance spectrum spans the wavelength range 3600–7400 Å, with gaps between 4820–4980 Å and 6160–6300 Å.
A.2.2 Nordic Optical Telescope
Long-slit spectroscopic observations were carried out at the Nordic Optical Telescope (NOT) using ALFOSC on 25 July 2025, at an airmass of approximately 1.5. A series of 13 180-second exposures were obtained using Grism #4, which covers the wavelength range from 0.32 µm to 0.96 µm, with a 1.8-arcsecond-wide slit. To minimize atmospheric dispersion, the instrument was aligned along the parallactic angle. No second-order blocking filter was used.
It should be noted that for Grism #4, second-order light from the -band becomes noticeable above 5900 Å, contributing up to 20% of the recorded signal, and from the -band contributes around 5%. For flux calibration and spectral normalization, the solar analogue BD-004074 and the spectrophotometric standard star Feige 110 were observed using the same setup.
We followed standard data reduction procedures. Helium and neon arc lamps were used for wavelength calibration, while halogen flats were employed for pixel response correction. Bias frames were also acquired to correct for the detector’s electronic offset.
Telescope | , km | , mag | , mag | , mag | , mag | , mag | , mag | ||
| Jul 2 | Lesedi | 10352 | 17.850.05 | 0.440.07 | 0.610.08 | 0.170.08 | – | – | ||
| Jul 3 | FTN | – | 18.080.02 | 0.640.02 | 0.990.03 | 0.350.02 | 0.230.04 | – | ||
| Jul 4 | FTS | – | 18.080.02 | 0.680.04 | 0.960.05 | 0.280.03 | 0.150.04 | – | ||
| Jul 5 | Skalnaté Pleso | 10418 | – | – | – | – | – | 0.380.02 | ||
| Jul 7 | Lesedi | 10634 | 17.310.07 | 0.670.09 | 0.810.09 | 0.140.10 | – | – | ||
| Jul 12 | FTN | – | 17.940.01 | 0.720.03 | 0.990.03 | 0.270.02 | 0.100.03 | – | ||
| Jul 13 | FTN | – | 17.860.02 | 0.610.05 | 0.900.05 | 0.290.04 | 0.100.04 | – | ||
| Jul 15 | Lesedi | 10661 | 17.200.05 | 0.690.08 | – | – | – | – | ||
| Jul 16 | FTN | – | 17.960.01 | 0.620.02 | 0.910.02 | 0.290.02 | 0.030.02 | – | ||
| Jul 18 | FTN/FTS | – | 17.800.02 | 0.620.04 | 0.880.04 | 0.270.04 | 0.100.04 | – | ||
| Jul 20 | Skalnaté Pleso | 11640 | – | – | – | – | – | 0.330.02 | ||
| Jul 21 | FTN | – | 17.780.02 | 0.670.03 | 0.960.02 | 0.290.03 | 0.080.02 | – | ||
| Jul 22 | Skalnaté Pleso | 11490 | – | – | – | – | – | 0.470.02 | ||
| Jul 23 | Lesedi | 10584 | 17.150.02 | 0.710.04 | – | – | – | – | ||
| Jul 23 | Skalnaté Pleso | 11418 | – | – | – | – | – | 0.510.02 | ||
| Jul 24 | FTS | – | 17.520.01 | 0.730.03 | 0.970.03 | 0.250.02 | 0.090.03 | – | ||
| Jul 26 | FTN | – | 17.710.01 | 0.640.02 | 0.940.02 | 0.300.02 | 0.100.03 | – | ||
| Jul 28 | FTN | – | 17.550.01 | 0.650.03 | 0.930.03 | 0.280.02 | 0.090.02 | – | ||
| Jul 29 | FTN | – | 17.730.01 | 0.670.02 | 0.940.02 | 0.270.02 | 0.060.02 | – | ||
| Average | – | – | – | 0.650.03 | 0.910.03 | 0.270.03 | 0.100.04 | 0.420.02 | ||
| ∗ The projected size of the photometric aperture at the distance of the comet. | ||||||||||
Appendix B Photometry and dust parameters derivation
| Jul 2 | Lesedi | – | – | – | – | – | 234 | ¡9 | |||||||||||||||||||||||||||||
| Jul 3 | Skalnaté Pleso | 10606 | 17.640.01 | 11.750.01 | – | – | – | – | 333 | ¡10 | |||||||||||||||||||||||||||
| Jul 5 | Skalnaté Pleso | 10418 | 17.470.01 | 11.650.01 | – | – | – | 375 | ¡11 | ||||||||||||||||||||||||||||
| Jul 7 | Lesedi | – | – | – | – | – | 308 | ¡10 | |||||||||||||||||||||||||||||
| Jul 15 | Lesedi | – | – | – | – | – | – | 237 | ¡9 | ||||||||||||||||||||||||||||
| Jul 20 | Skalnaté Pleso | 11640 | 17.330.01 | 12.010.01 | – | – | – | 329 | ¡10 | ||||||||||||||||||||||||||||
| Jul 22 | Skalnaté Pleso | 11490 | 17.220.01 | 11.970.01 | – | – | – | 355 | ¡11 | ||||||||||||||||||||||||||||
| Jul 23 | Lesedi | – | – | – | – | – | – | 431 | ¡11 | ||||||||||||||||||||||||||||
| Jul 23 | Skalnaté Pleso | – | – | – | – | – | – | 359 | ¡11 | ||||||||||||||||||||||||||||
| Average | – | – | – | – | 29811 | 35113 | 34518 | 28126 | 29422 | 329 | ¡10 | ||||||||||||||||||||||||||
| ∗ ∗ - the projected size of the photometric aperture at the distance of the comet. | |||||||||||||||||||||||||||||||||||||
The comet magnitudes in various filters on 02, 07, and 15 July 2025, measured within an aperture ranging from 10,300 to 11,600 km in radius centered on the optocentre, are listed in Table 1. Standard star magnitudes were taken from the APASS catalog (Henden et al. 2016), with an average photometric uncertainty of 0.02 mag.
Absolute magnitudes were computed as:
where and are the heliocentric and geocentric distances (in au), is the phase angle, and mag/deg is the assumed phase coefficient.
The dust production proxy was calculated following A’Hearn et al. (1984):
where and are the magnitudes of the Sun and the comet, and is the aperture radius. Solar magnitudes were taken from Willmer (2018). Phase corrections to phase angle were applied using the function from Schleicher & Bair (2011). Results are provided in Tables 1 and 2.
The scattering cross-section (km2) was derived following Jewitt et al. (2016):
where is the geometric albedo. The corresponding radius is , representing an upper limit to the nucleus size.
Dust ejection velocity estimates follow Jewitt & Meech (1987) and Hsieh et al. (2021), yielding velocities between 0.012 and 1.2 m s-1 for particle sizes from millimeters to meters.
The dust mass loss rate was calculated following Fink & Rubin (2012):
Appendix C Determination of the spin state and lightcurve amplitude
We compiled all observations obtained in the SDSS band to derive the lightcurve of 3I/ATLAS. For some datasets, particularly those obtained during the first week of July, image stacking was employed to increase the signal-to-noise ratio (SNR) of the comet. All individual frames were visually inspected to identify and exclude those affected by stellar contamination, which was common due to the comet’s motion toward the Galactic plane. To maximize the signal from the cometary nucleus while minimizing the contribution from the surrounding coma and background noise, photometric measurements were performed using an optimized circular aperture matched to the object’s point spread function (PSF). Specifically, the aperture radius was set to encompass approximately one standard deviation () of the full width at half maximum (FWHM) of the comet’s brightness profile in the stacked image. Photometric uncertainties were computed by combining photon noise, background noise, and scatter in the reference star magnitudes.
The photometric measurements were reduced by applying light-time correction to the observation epochs, as well as distance corrections for the varying heliocentric and geocentric distances. A phase correction was applied using the H–G photometric system (Bowell et al. 1989), assuming a standard slope parameter of . To determine the best-fitting rotation period, we performed a Fourier analysis using a second-order harmonic function (see Fig 6).